OpenStudy (anonymous):

x, y, z is a geometric sequence with common ratio r and \[x \neq y\] . if x, 2y, 3z is an arithmetic sequence, find r.

OpenStudy (unklerhaukus):

is that the whole question?

OpenStudy (anonymous):

yes... what's r in the geometric sequence..

OpenStudy (unklerhaukus):

y/x ?

OpenStudy (anonymous):

I think \[r=\lim_{n \rightarrow \infty}x _{n}/y _{n}\]

OpenStudy (unklerhaukus):

so, 1/3?

OpenStudy (anonymous):

yeah.. @UnkleRhaukus .... good one...

OpenStudy (unklerhaukus):

\[3\approx\infty\]

OpenStudy (anonymous):

are you sure this is the answer?

OpenStudy (unklerhaukus):

no

OpenStudy (anonymous):

i don't know, but i see it this way: y=rx,z=r^2x so: 2y-x=q or 2rx-x=q also: 3r^2x-x=2q substituting value for q in the 2º equation: 3r^2x-x=4rx-2x or: r^2-4/3r-1=0 and solve for r

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

is the answer 1/3 in question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have no idea how he got 1/3

OpenStudy (anonymous):

y=xr, z=yr=xr^2 3z - 2y = 2y - x

OpenStudy (anonymous):

substitute y an z into that last equation..

OpenStudy (anonymous):

oh, right, i made a mistake....

OpenStudy (anonymous):

you were getting it though....

OpenStudy (anonymous):

...